log8^(n-3)+log8(n+4)=1
Answer by stanbon(75887)

log8(n-3)+log8(n+4)=1
----
log8[n^2+n-12] = 1
----
n^2+n-12 = 8
n^2+n-20 = 0
(n+5)(n-4) = 0
---
n cannot be -5 because log8(-8) is meaningless.
---
n = 4 gives log8(4-3)+log8(4+4) = 1
gives log8(1) + log8(8) = 1
so 0+1 = 1
1 = 1
----
Solution: n=4
============================
Cheers,
Stan H.